Brown will receive $1,000 at the end of 10 years. In return, he will pay $200 now and another payment in 5 years. At a 9% discount rate, what should the size of Browns second payment be?
Answer #1

PV of Payments paid = PV of Amount received

200/(1+9%) + Second Payment/(1+9%)5 = 1000/(1+9%)10
Second Payment/1.095 = 1000/1.0910 - 200/1.09
Second payment/1.095 = 238.9246
Second Payment = 238.9246 * 1.095 = 367.62

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well